# NeuroBench: An Open-Source Benchmark Framework for the Standardization of Methodology in Brainwave-based Authentication Research

### 14 Feb 2024 v2

@@ Abstract
Biometric systems based on brain activity have been proposed as an alternative to passwords or to complement current authentication techniques. By leveraging the unique brainwave patterns of individuals, these systems offer the possibility of creating authentication solutions that are resistant to theft, hands-free, accessible, and potentially even revocable. However, despite the growing stream of research in this area, faster advance is hindered by reproducibility problems. Issues such as the lack of standard reporting schemes for performance results and system configuration, or the absence of common evaluation benchmarks, make comparability and proper assessment of different biometric solutions challenging. Further, barriers are erected to future work when, as so often, source code is not published open access. To bridge this gap, we introduce NeuroBench, a flexible open source tool to benchmark brainwave-based authentication models. It incorporates nine diverse datasets, implements a comprehensive set of pre-processing parameters and machine learning algorithms, enables testing under two common adversary models (known vs unknown attacker), and allows researchers to generate full performance reports and visualizations. We use NeuroBench to investigate the shallow classifiers and deep learning-based approaches proposed in the literature, and to test robustness across multiple sessions. We observe a 37.6% reduction in Equal Error Rate (EER) for unknown attacker scenarios (typically not tested in the literature), and we highlight the importance of session variability to brainwave authentication. All in all, our results demonstrate the viability and relevance of NeuroBench in streamlining fair comparisons of algorithms, thereby furthering the advancement of brainwave-based authentication through robust methodological practices.
@@ 1 Introduction
  • (1) We present NeuroBench, the first comprehensive open-source [^1] https://github.com/Avichaurasia/Brain-Models benchmark tool designed to assist researchers in evaluating their brainwave-based authentication approaches. NeuroBench is engineered to allow for the easy and flexible integration of new datasets and methods, enabling fair comparability. NeuroBench incorporates nine public datasets, which together contain a substantial number of subjects (n = 285, Table 2) and multi-session recordings. To determine the impact of preprocessing on performance, NeuroBench allows to apply different sample rejection thresholds as well as sample length parameters (Sec 4.1) and feature extraction methods (Sec 4.2). Also, it implements a baseline with the most popular authentication algorithms in the state of the art, including shallow classifiers and deep learning-based approaches. Lastly, it incorporates two attack scenarios under which to test the algorithms mansfield2002best : i) known attacker, where the attacker is known to the authentication model (i.e., enrolled in the system), and, ii) unknown attacker, a more realistic situation in which the attacker biometric data has not been seen by the system previous to the attack.
  • (2) We use NeuroBench to build a comprehensive benchmark of popular brainwave-based authentication algorithms, providing the first unbiased comparison of approaches. Namely, we compare six shallow classifiers: SVM (Support Vector Machine), Random Forest, KNN (K-Nearest Neighbors), LDA (Linear Discriminant Analysis), Naive Bayes (NB), and Logistic Regression (LR). Besides this benchmark, which covers the biggest share of solutions in the literature, we provide a second benchmark for the recent stream of work on deep learning-based approaches. Specifically, NeuroBench implements the representative similarity-based technique using Twin Neural Networks (Sec. 4.5). Beyond performance across an extensive list of publicly available EEG datasets, we focus our comparative analysis on two under-explored aspects in current research: How robust are different solutions in multi-session scenarios? (Sec. 4.4) and What is the impact of the adversary model? (Sec. 4.3)
@@ 2.1 Using brainwaves for Biometric Authentication
    Raw EEG Data Acquisition: This initial step captures the brain’s electrical activity. An effective way to elicit signals useful for authentication is measuring brain reactions to controlled stimuli, such as audio or images. These time-locked reactions,called Event-Related Potentials (ERPs), have shown unique features to individuate people arias2021inexpensive and are distinguished by their high Signal-to-Noise Ratio (SNR) armstrong2015brainprint ; zhang2021review . Moreover, the ability to modify the stimuli associated with ERPs enhances the system’s revocability. This feature is invaluable in mitigating risks associated with potential data compromise lin2018brain . NeuroBench focuses in ERP-based datasets, given their superior benefits with regard to alternative brainwave elicitation paradigms.
@@
    Signal Pre-processing: The acquired raw EEG data is then refined through pre-processing. This may include filtering out noise, rejecting artifacts, and downsampling the signal, aiming to isolate the purest form of the EEG data for authentication. purposes gui2019survey . Neurobench embeds a pre-processing pipeline based on best-practices for handling ERPs as used in the community gui2019survey .
@@
    Feature Extraction: The pre-processed EEG signals are further analyzed to extract key features. These features can be directly obtained from the ERP EEG signal, such as power spectral densities, or generated as representation vectors learned by a neural network. Neurobench implements both kind of approaches.
@@
    Decision Maker: In the authentication phase, the system evaluates the incoming EEG sample, against the stored model/template, applying similarity thresholds to decide if it belongs to a legitimate user or to an impostor. NeuroBench incorporates the most common techniques for machine learning model-based approaches (SVM, KNN, LDA, NB, LR, RF), as well as the main approach for template-based authentication using deep neural networks (Twin Neural Networks).
@@ 2.2 Adversary Model
Figure 2: Schematic depiction outlining the architectural structure of the NeuroBench codebase, providing a visual overview of its underlying framework jayaram2018moabb
@@ 2.3 Related Work
3 NeuroBench
As illustrated in Figure 2, NeuroBench is organized into five integral components: dataset, preprocessing, feature extraction, evaluation, and analysis. Our benchmarking workflow is influenced by MOABB (Mother of all BCI benchmarks) jayaram2018moabb work. In line with our research aims, we have tailored and refined their methodology to create a customized benchmarking suite specifically designed for evaluating and analyzing brainwave authentication systems. In the following, we detail NeuroBench’s components, justifying the design choices.
@@ 3.3 Feature Extraction
Furthermore, NeuroBench implements a deep learning based approach using a TNN with triplet loss function to transform time-series EEG data into compact brain embeddings. Building on the work of Fallahi et al. fallahi2023brainnet , our TNN structure integrates three Convolutional Neural Network (CNN) branches, each featuring five convolutional layers, as illustrated in Figure 3. These branches received 1-second samples as input, structured in a two-dimensional array where rows correspond to channel indices, and columns represent discrete-time measurements. After each convolutional layer, we applied an average pooling layer to reduce the input vectors’ dimensionality while preserving the unique characteristics of each brainwave.
@@ 4.6 Comparison with Related Work
The comparison of our results will initially focus on the results published by Arias et al. arias2021inexpensive and Fallahi et al. fallahi2023brainnet . This is due to the establishment of our benchmark, which is based on their implementations. Moreover, both studies utilized the ERPCORE erpcore P300 and N400 datasets, similar to our approach. Arias et al. arias2021inexpensive investigate the performance of different shallow classifiers within single-session brainwave authentication. The results indicate 1.9% EER for N400 and 3% EER for P300 paradigm with a sample rejection of 120 µV and RF classifier, we observe the same results where we achieved 2.9% EER for P300 and 1.13% EER to N400 paradigm with the sample rejection of 100 µV and RF classifier. Also, Fallahi et al. fallahi2023brainnet acquired 2.01% and 1.37% respectively for P300 and N400 ERPCORE and 0.14 for P300:bi2015a dataset bi2015a , where we obtained %1.53, 0.83% , and 0.43 respectively. As can be observed the results are fairly similar and trends are the same. The small difference in results could be from a different random seed, version of libraries, and some evaluation parameters like epoch rejection rate for shallow classifiers and number of epochs for twin neural networks. To enhance the reproducibility of our result, we provided a docker container in the GitHub of NeuroBench.
@@ 5 Conclusion and Future Work
In conclusion, NeuroBench is presented as a comprehensive benchmarking tool for EEG authentication, particularly supporting ERP-based methods. As an open-source platform, it currently supports 9 datasets and 7 classifiers. It can easily be adapted to include new datasets and classifiers, allowing for the investigation of different methodological research questions. Our results highlight that epoch rejection does not significantly affect outcomes. PSD combined with AR order 1 emerges as a recommended default feature set for shallow classifiers. Known attacker evaluation can misleadingly indicate lower EER. Multi-session evaluation is significantly more challenging than single-session evaluation. Moreover, it appears that due to the limited availability of data, deep learning approaches do not consistently outperform shallow classifiers in brainwave-based authentication.
@@ Appendix B Installation and Running the tool
The benchmarking tool is available on our Github repository  [^4] https://github.com/Avichaurasia/Brain-Models. Developed in Python, the tool leverages various statistical and machine-learning libraries, necessitating the establishment of a Python environment as a preliminary step. The subsequent procedures delineate the steps to set up the operational envirmoment and replicate outcomes utilizing the NeuroBench tool.

### 07 Apr 2024 v3

# NeuroIDBench: An Open-Source Benchmark Framework for the Standardization of Methodology in Brainwave-based Authentication Research

@@ Abstract
Biometric systems based on brain activity have been proposed as an alternative to passwords or to complement current authentication techniques. By leveraging the unique brainwave patterns of individuals, these systems offer the possibility of creating authentication solutions that are resistant to theft, hands-free, accessible, and potentially even revocable. However, despite the growing stream of research in this area, faster advance is hindered by reproducibility problems. Issues such as the lack of standard reporting schemes for performance results and system configuration, or the absence of common evaluation benchmarks, make comparability and proper assessment of different biometric solutions challenging. Further, barriers are erected to future work when, as so often, source code is not published open access. To bridge this gap, we introduce NeuroIDBench, a flexible open source tool to benchmark brainwave-based authentication models. It incorporates nine diverse datasets, implements a comprehensive set of pre-processing parameters and machine learning algorithms, enables testing under two common adversary models (known vs unknown attacker), and allows researchers to generate full performance reports and visualizations. We use NeuroIDBench to investigate the shallow classifiers and deep learning-based approaches proposed in the literature, and to test robustness across multiple sessions. We observe a 37.6% reduction in Equal Error Rate (EER) for unknown attacker scenarios (typically not tested in the literature), and we highlight the importance of session variability to brainwave authentication. All in all, our results demonstrate the viability and relevance of NeuroIDBench in streamlining fair comparisons of algorithms, thereby furthering the advancement of brainwave-based authentication through robust methodological practices.
@@ 1 Introduction
  • (1) We present NeuroIDBench, the first comprehensive open-source [^1] https://github.com/Avichaurasia/NeuroIDBench.git benchmark tool designed to assist researchers in evaluating their brainwave-based authentication approaches. NeuroIDBench is engineered to allow for the easy and flexible integration of new datasets and methods, enabling fair comparability. NeuroIDBench incorporates nine public datasets, which together contain a substantial number of subjects (n = 285, Table 2) and multi-session recordings. To determine the impact of preprocessing on performance, NeuroIDBench allows to apply different sample rejection thresholds as well as sample length parameters (Sec 4.1) and feature extraction methods (Sec 4.2). Also, it implements a baseline with the most popular authentication algorithms in the state of the art, including shallow classifiers and deep learning-based approaches. Lastly, it incorporates two attack scenarios under which to test the algorithms mansfield2002best : i) known attacker, where the attacker is known to the authentication model (i.e., enrolled in the system), and, ii) unknown attacker, a more realistic situation in which the attacker biometric data has not been seen by the system previous to the attack.
  • (2) We use NeuroIDBench to build a comprehensive benchmark of popular brainwave-based authentication algorithms, providing the first unbiased comparison of approaches. Namely, we compare six shallow classifiers: SVM (Support Vector Machine), Random Forest, KNN (K-Nearest Neighbors), LDA (Linear Discriminant Analysis), Naive Bayes (NB), and Logistic Regression (LR). Besides this benchmark, which covers the biggest share of solutions in the literature, we provide a second benchmark for the recent stream of work on deep learning-based approaches. Specifically, NeuroIDBench implements the representative similarity-based technique using Twin Neural Networks (Sec. 4.5). Beyond performance across an extensive list of publicly available EEG datasets, we focus our comparative analysis on two under-explored aspects in current research: How robust are different solutions in multi-session scenarios? (Sec. 4.4) and What is the impact of the adversary model? (Sec. 4.3)
@@ 2.1 Using brainwaves for Biometric Authentication
    Raw EEG Data Acquisition: This initial step captures the brain’s electrical activity. An effective way to elicit signals useful for authentication is measuring brain reactions to controlled stimuli, such as audio or images. These time-locked reactions,called Event-Related Potentials (ERPs), have shown unique features to individuate people arias2021inexpensive and are distinguished by their high Signal-to-Noise Ratio (SNR) armstrong2015brainprint ; zhang2021review . Moreover, the ability to modify the stimuli associated with ERPs enhances the system’s revocability. This feature is invaluable in mitigating risks associated with potential data compromise lin2018brain . NeuroIDBench focuses in ERP-based datasets, given their superior benefits with regard to alternative brainwave elicitation paradigms.
@@
    Signal Pre-processing: The acquired raw EEG data is then refined through pre-processing. This may include filtering out noise, rejecting artifacts, and downsampling the signal, aiming to isolate the purest form of the EEG data for authentication. purposes gui2019survey . NeuroIDBench embeds a pre-processing pipeline based on best-practices for handling ERPs as used in the community gui2019survey .
@@
    Feature Extraction: The pre-processed EEG signals are further analyzed to extract key features. These features can be directly obtained from the ERP EEG signal, such as power spectral densities, or generated as representation vectors learned by a neural network. NeuroIDBench implements both kind of approaches.
@@
    Decision Maker: In the authentication phase, the system evaluates the incoming EEG sample, against the stored model/template, applying similarity thresholds to decide if it belongs to a legitimate user or to an impostor. NeuroIDBench incorporates the most common techniques for machine learning model-based approaches (SVM, KNN, LDA, NB, LR, RF), as well as the main approach for template-based authentication using deep neural networks (Twin Neural Networks).
@@ 2.2 Adversary Model
Figure 2: Schematic depiction outlining the architectural structure of the NeuroIDBench codebase, providing a visual overview of its underlying framework jayaram2018moabb
@@ 2.3 Related Work
3 NeuroIDBench
As illustrated in Figure 2, NeuroIDBench is organized into five integral components: dataset, preprocessing, feature extraction, evaluation, and analysis. Our benchmarking workflow is influenced by MOABB (Mother of all BCI benchmarks) jayaram2018moabb work. In line with our research aims, we have tailored and refined their methodology to create a customized benchmarking suite specifically designed for evaluating and analyzing brainwave authentication systems. In the following, we detail NeuroIDBench’s components, justifying the design choices.
@@ 3.3 Feature Extraction
Furthermore, NeuroIDBench implements a deep learning based approach using a TNN with triplet loss function to transform time-series EEG data into compact brain embeddings. Building on the work of Fallahi et al. fallahi2023brainnet , our TNN structure integrates three Convolutional Neural Network (CNN) branches, each featuring five convolutional layers, as illustrated in Figure 3. These branches received 1-second samples as input, structured in a two-dimensional array where rows correspond to channel indices, and columns represent discrete-time measurements. After each convolutional layer, we applied an average pooling layer to reduce the input vectors’ dimensionality while preserving the unique characteristics of each brainwave.
@@ 4.6 Comparison with Related Work
The comparison of our results will initially focus on the results published by Arias et al. arias2021inexpensive and Fallahi et al. fallahi2023brainnet . This is due to the establishment of our benchmark, which is based on their implementations. Moreover, both studies utilized the ERPCORE erpcore P300 and N400 datasets, similar to our approach. Arias et al. arias2021inexpensive investigate the performance of different shallow classifiers within single-session brainwave authentication. The results indicate 1.9% EER for N400 and 3% EER for P300 paradigm with a sample rejection of 120 µV and RF classifier, we observe the same results where we achieved 2.9% EER for P300 and 1.13% EER to N400 paradigm with the sample rejection of 100 µV and RF classifier. Also, Fallahi et al. fallahi2023brainnet acquired 2.01% and 1.37% respectively for P300 and N400 ERPCORE and 0.14 for P300:bi2015a dataset bi2015a , where we obtained %1.53, 0.83% , and 0.43 respectively. As can be observed the results are fairly similar and trends are the same. The small difference in results could be from a different random seed, version of libraries, and some evaluation parameters like epoch rejection rate for shallow classifiers and number of epochs for twin neural networks. To enhance the reproducibility of our result, we provided a docker container in the GitHub of NeuroIDBench.
@@ 5 Conclusion and Future Work
In conclusion, NeuroIDBench is presented as a comprehensive benchmarking tool for EEG authentication, particularly supporting ERP-based methods. As an open-source platform, it currently supports 9 datasets and 7 classifiers. It can easily be adapted to include new datasets and classifiers, allowing for the investigation of different methodological research questions. Our results highlight that epoch rejection does not significantly affect outcomes. PSD combined with AR order 1 emerges as a recommended default feature set for shallow classifiers. Known attacker evaluation can misleadingly indicate lower EER. Multi-session evaluation is significantly more challenging than single-session evaluation. Moreover, it appears that due to the limited availability of data, deep learning approaches do not consistently outperform shallow classifiers in brainwave-based authentication.
@@ Appendix A Receiver Operating Characteristic (ROC) Curves
ROC curves play a pivotal role in assessing the performance of biometric systems like brainwave authentication. These curves provide a graphical representation of the system’s ability to distinguish between genuine users and impostors by plotting FMR against 1-FNMR across various threshold values. In brainwave authentication systems, ROC curves help in determining the optimal threshold for achieving the desired balance between sensitivity and specificity, thereby ensuring the system’s effectiveness in accurately identifying authorized users while minimizing the risk of false positives and false negatives.
@@ Appendix B Installation and Running the tool
The benchmarking tool is available on our Github repository  [^4] https://github.com/Avichaurasia/NeuroIDBench.git. Developed in Python, the tool leverages various statistical and machine-learning libraries, necessitating the establishment of a Python environment as a preliminary step. The subsequent procedures delineate the steps to set up the operational envirmoment and replicate outcomes utilizing the NeuroIDBench tool.